# Powerformer: A Transformer with Weighted Causal Attention for Time-series Forecasting

### v1

@@ Acknowledgements
Powerformer forecasts univariate time-series by combining WCMHA, the commonly used encoder-only Transformer architecture, and patching [27]. We selected PatchTST as the base of Powerformer since PatchTST already includes decomposes multivariate time-series into univariate channels and performs the patching. Additionally, similar to other modern time-series Transformer-based models [49, 24, 32], PatchTST [27] uses an encoder-only architecture with a linear readout head. Powerformer uses a standard Transformer encoder architecture with the primary difference being the replacement of the vanilla MHA by weighted causal WCMHA, described in Eqs. 6-10. We additionally remove the dropout applied to the attention weights because our implicit biases enforce a deterministic temporal dependency on pairwise correlations. Dropout enforces redundancy between pairwise correlations and is therefore in conflict with our implicit biases. The input data comes in with $D$ variates and shape $[B,D,T_{\text{seq}}]$, where $B$ is the batch size and $T_{\text{seq}}$ is the input sequence length. The PatchTST patching mechanism (described in Appendix A3.2) returns the patched and normalized univariate embeddings with shape $[B\times D,P,N]$, where $P$ is the number of patches and $N$ is the size of the embeddings. These embeddings are encoded through multiple WCMHA Transformer encoder layers (see Appendix A3.4 for hyperparameter details). The encoder output (of shape $[B\times D,P,N]$) is flattened along the last two dimensions ($[B\times D,P\times N]$) in preparation for the linear readout head. The linear readout matrix has the shape $[P\times N,T_{\text{pred}}]$ and after acting on the flattened encoder output produces the univariate forecasts of shape $[B\times D,T_{\text{pred}}]$. Finally, these univariate forecasts are renormalized and reshaped into the multivariate input structure $[B,D,T_{\text{pred}}]$ and returned to the user. Consequently, Powerformer is a simple Transformer-based model with causal and local masks added to the attention scores $\mathbf{S}$, and patching applied to the input data. We present the Powerformer hyperparameters used in this work in Table A2.
A3.4 Model Tuning Parameters
@@
To reproduce our results, we list our hyperparameters used on each dataset for Powerformer (Table A2) and Transformer (Table A3). For each dataset and model, we trained with the following 3 random seeds: 2021, 1776, and 1953.
@@
Table A2: We provide Powerformer architecture and training parameters used for each dataset.
@@
Learning Rate $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$
@@
Table A3: We provide Transformer architecture and training parameters used for each dataset.
@@
Here, we show Powerformer forecasting visualizations and compare various decay time constants against ground truth and PatcTST [27]. We also highlight plots where both Powerformer and PatchTST struggle. In these cases Powerformer generally outperforms better capture faster-moving features, non-periodic features, and is better able to alter periodic predictions. We believe this is due to Powerformer’s ability to focus on and amplify local high-frequency features while dampening or removing long-time high-frequency components. WCMHA achieves this through the power-law masks $f^{(\text{PL})}(t)$ and $f^{(\text{SPL})}(t)$, which amplify local correlations and dampen long-time correlations.
@@
Figure A6: We show forecasting results on the Electricity dataset for Powerformer with $f^{(\text{PL})}(t)$ and PatchTST. The colored lines represent PatchTST and different $f^{(\text{PL})}(t)$ decay constants. For these forecasts, the sequence length is 512 and the prediction length is 336.
@@
Figure A7: We show forecasting results on the ETTm1 dataset for Powerformer with $f^{(\text{PL})}(t)$ and PatchTST. The colored lines represent PatchTST and different $f^{(\text{PL})}(t)$ decay constants. For these forecasts, the sequence length is 512 and the prediction length is 336.
@@
Figure A8: We show forecasting results on the Weather dataset for Powerformer with $f^{(\text{PL})}(t)$ and PatchTST. The colored lines represent PatchTST and different $f^{(\text{PL})}(t)$ decay constants. For these forecasts, the sequence length is 512 and the prediction length is 336.
Appendix A5 Experiments
A5.1 Experiment Details
@@
All experiments are evaluated 3 times with different initialization (A3.4). For each dataset, we select a single input sequence based on the best-performing input sequence length across all prediction lengths. We choose a single input sequence length for each dataset because the strength of the correlations between time steps is independent of the prediction length, but the required input sequence length to capture the information necessary to forecast is unique for each dataset. For each prediction length, we treat the attention mask as a hyperparameter and select the best mask and time constant pair. We select attention masks for each prediction length because the forecasting timescale will likely emphasize short or long-term correlations for shorter or longer forecasting windows. We make our selections based on the MSE error.
A5.2 Powerformer
@@
We evaluate the weight power-law ($f^{(\text{PL})}(t)$) and similarity power-law ($f^{(\text{SPL})}(t)$) filters on all 7 datasets. For $f^{(\text{PL})}(t)$ we evaluate $\alpha\in\{0.1,0.25,0.5,0.75,1.0\}$ and for $f^{(\text{SPL})}(t)$ we evaluate $\alpha\in\{0.1,0.5,1.0,2.0\}$. Similar to PatchTST [27] we evaluate the power-law filters with two input lengths: 336 and 512.
@@
We train a standard encoder-decoder Transformer architecture both with and without WCMHA. Generic Transformer models often incorporate causality through masking. To account for this, our Transformer benchmark uses masking for the decoder MHA but not for the encoder MHA nor the decoder cross-attention. We investigate WCMHA’s effects by replacing the encoder and decoder self-attention, but not the decoder cross-attention as it is already causal. Section A3.1 provides more architectural details.
@@
We train Transformer models, with and without WCMHA, using the weight power-law ($f^{(\text{PL})}(t)$) filter on all 7 datasets. We explore the following decay constants: $\alpha\in[0.1,0.25,0.5,0.75,1.0]$ and provide further experimental detail in sections 4 and A5.1.
A5.4 Baseline Selection
@@
We compare Powerformer against multiple Transformer-based state-of-the-art methods. We populate Table 1 with values taken from the models’ respective papers. All models except TOTEM [32] and iTransformer [24] treat the input length as a hyperparameter and report the best results. PatchTST [27] provides results for two input sequence lengths (336 and 512). In Table 1 we report the best result between these input sequence lengths for each dataset based on the MSE. We employ the same input sequence length selection for Powerformer.
Appendix A6 Ablation Studies
Here, we perform ablation experiments on WCMHA in both Transformer and Powerformer. For both models, we examine the effects of different masks ($f^{(\text{PL})}(t)$, $f^{(\text{SPL})}(t)$, and $f^{(\text{BW})}_{n}(t)$) and various time decays $\alpha$. We do not perform the same amount of experiments for each model and mask pairing due to poor performance. In the remainder of this section, we provide detailed tables of the MSE and MAE scores, as well as plots of the attention scores and weights. In the last subsection (A6.3), we investigate the effects of making $\alpha$ a learnable parameter.
A6.1 Transformer Results
@@
Here, we present the full results from evaluating Transformer with WCMHA. The architecture is described in Appendix A3.1 and the experimental details are provided in Appendices A5.1 and A5.3.
@@
Table A4 provides the aggregate Transformer performance without WCMHA and with WCMHA for varying decay times ($\alpha$). WCMHA decisively outperforms MHA on 4 or the 7 datasets and ties with the number of best results on ETTh2. Looking further into ETTh2 we see that when WCMHA outperforms MHA it does so by much larger margins (16%) than when it underperforms (2%). Moreover, we observe the same larger margins of overperformance across all of the datasets. This indicates that when the natural pairwise distribution is found we can see significant improvements in MSE and MAE, but the effects of imposing the wrong pairwise distribution are much less significant.
@@
In Figs. A9-A22 we present the attention score and weight distributions with and without applying the local-causal mask. The distributions can have significant deviations between datasets. In general, the WCMHA decoder self-attention shows the smallest deviation from the MHA base case. We note that the MHA base case already has a causal mask. The last WCMHA encoder self-attention layer often shows the largest difference between WCMHA and MHA distributions. This makes sense since the MHA encoder attention is not causal. We do not alter the decoder cross-attention, but we still observe shifts in the distribution, which differ between datasets. Interestingly we do not observe the same bimodal distribution in attention weights as we do for Powerformer. This may be due to the Transformer’s encoder-decoder structure, whereas Powerformer only has an encoder.
@@
Table A4: We provide the forecasting MSE and MAE on the test sets for Transformer with MHA (base case) and for Transformer with WCMHA and the weight power-law mask $f^{(\text{PL})}(t)$. The top row indicates the model or decay constant $(\alpha)$, the bold numbers are the best (lowest) performance, and the underlined numbers are the second best.
@@
Figure A9: We present the Transformer attention score (inset) and weight distributions on the ETTh1 dataset with a forecasting length of 96. The dotted line represents the reference MHA Transformer results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A10: We present the Transformer attention score (inset) and weight distributions on the ETTh1 dataset with a forecasting length of 720. The dotted line represents the reference MHA Transformer results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A11: We present the Transformer attention score (inset) and weight distributions on the ETTh2 dataset with a forecasting length of 96. The dotted line represents the reference MHA Transformer results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A12: We present the Transformer attention score (inset) and weight distributions on the ETTh2 dataset with a forecasting length of 720. The dotted line represents the reference MHA Transformer results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A13: We present the Transformer attention score (inset) and weight distributions on the ETTm1 dataset with a forecasting length of 96. The dotted line represents the reference MHA Transformer results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A14: We present the Transformer attention score (inset) and weight distributions on the ETTm1 dataset with a forecasting length of 720. The dotted line represents the reference MHA Transformer results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A15: We present the Transformer attention score (inset) and weight distributions on the ETTm2 dataset with a forecasting length of 96. The dotted line represents the reference MHA Transformer results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A16: We present the Transformer attention score (inset) and weight distributions on the ETTm2 dataset with a forecasting length of 720. The dotted line represents the reference MHA Transformer results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
@@
Figure A17: We present the Transformer attention score (inset) and weight distributions on the Weather dataset with a forecasting length of 96. The dotted line represents the reference MHA Transformer results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
@@
Figure A18: We present the Transformer attention score (inset) and weight distributions on the Weather dataset with a forecasting length of 720. The dotted line represents the reference MHA Transformer results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
@@
Figure A19: We present the Transformer attention score (inset) and weight distributions on the Electricity dataset with a forecasting length of 96. The dotted line represents the reference MHA Transformer results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
@@
Figure A20: We present the Transformer attention score (inset) and weight distributions on the Electricity dataset with a forecasting length of 720. The dotted line represents the reference MHA Transformer results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A21: We present the Transformer attention score (inset) and weight distributions on the Traffic dataset with a forecasting length of 96. The dotted line represents the reference MHA Transformer results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A22: We present the Transformer attention score (inset) and weight distributions on the Traffic dataset with a forecasting length of 720. The dotted line represents the reference MHA Transformer results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
A6.2 Powerformer
@@
We present the full Powerformer results on all the evaluated masks: $f^{(\text{PL})}(t)$, $f^{(\text{SPL})}(t)$, $f^{(\text{BW})}_{1}(t)$, and $f^{(\text{BW})}_{2}(t)$. Architecture details can be found in Appendix A3.3, while Appendices A5.1 and A5.2 outline the experimental details.
A6.2.1 Weight Power-Law Mask
@@
Table A5 presents the aggregate MSE and MAE results for weight power-law mask ($f^{(\text{PL})}(t)$) with varying time decays ($\alpha$), forecasting lengths, and input sequence lengths. The 512 input sequence generally outperforms the shorter 336 input length. As expected, the best performing $\alpha$ varies between datasets, but it also varies as a function of forecast length and input sequence length. Given that some of the best-performing decay lengths are $\alpha=1$ with further experiments one may improve upon these values. However, our $f^{(\text{SPL})}(t)$ mask includes faster-decaying weights, as shown in Fig. 2.
@@
We compare all datasets’ attention score and weight distributions both with MHA and WCMHA in Figs A23-A36. We observe similar bimodal distributions for all datasets and note wider $\mathbf{C}^{(\text{C,D})}$ distributions for ETTm1, ETTm2, and Weather.
@@
Table A5: We compare Powerformer’s performance with the weight power law mask $f^{(\text{PL})}(t)$ on standard time-series datasets for varying decay lengths. The best results are bolded and second best are underlined.

### v2

@@ Acknowledgements
Powerformer forecasts univariate time-series by combining RBCA, the commonly used encoder-only Transformer architecture, and patching (Nie et al., 2023). We selected PatchTST as the base of Powerformer since PatchTST already includes decomposes multivariate time-series into univariate channels and performs the patching. Additionally, similar to other modern time-series Transformer-based models (Zhou et al., 2023; Liu et al., 2024; Talukder et al., 2024), PatchTST (Nie et al., 2023) uses an encoder-only architecture with a linear readout head. Powerformer uses a standard Transformer encoder architecture with the primary difference being the replacement of the vanilla MHA by RBCA. We additionally remove the dropout applied to the attention weights because our implicit biases enforce a deterministic temporal dependency on pairwise correlations. Dropout enforces redundancy between pairwise correlations and is therefore in conflict with our implicit biases. The input data comes in with $D$ variates and shape $B,D,T_{\text{seq}}]$, where $B$ is the batch size and $T_{\text{seq}}$ is the input sequence length. The PatchTST patching mechanism (described in Appendix S3.2) returns the patched and normalized univariate embeddings with shape $B\times D,P,N]$, where $P$ is the number of patches and $N$ is the size of the embeddings. These embeddings are encoded through multiple RBCA Transformer encoder layers (see Appendix S3.4 for hyperparameter details). The encoder output (of shape $B\times D,P,N]$) is flattened along the last two dimensions ($B\times D,P\times N]$) in preparation for the linear readout head. The linear readout matrix has the shape $P\times N,T_{\text{pred}}]$ and after acting on the flattened encoder output produces the univariate forecasts of shape $B\times D,T_{\text{pred}}]$. Finally, these univariate forecasts are renormalized and reshaped into the multivariate input structure $B,D,T_{\text{pred}}]$ and returned to the user. Consequently, Powerformer is a simple Transformer-based model with causal and local masks added to the attention scores $\mathbf{S}$, and patching applied to the input data. We present the Powerformer hyperparameters used in this work in Table S2.
S3.4 Hyperparameter Tuning and Selection
In addition to tuning standard Transformer hyperparameters, Powerformer introduces two patching and two recency-bias hyperparameters. The patching mechanism is determined by the patch length and the patch stride. The patch length is the number of input tokens that are averaged together. The patch stride is the stride of the patching mechanism that helps reduce the size of the input data. For the recency-bias mask, one can tune the type of mask functional and it’s decay time. As previously described, and illustrated in Fig. 2, we test four masks: power-law mask ($f^{(\text{PL})}(t)$), similarity power-law mask ($f^{(\text{SPL})}(t)$), and two Butterworth masks for $n\in{1,2}$ ($f^{(\text{BW})}_{n}(t)$). For each mask, one may try different decay constants $\alpha$ that determine how fast the mask decays. In this works we tried the following decay constants for each mask: for $f^{(\text{PL})}(t)$ we evaluate $\alpha\in\{0.1,0.25,0.5,0.75,1.0\}$; for $f^{(\text{SPL})}(t)$ we evaluate $\alpha\in\{0.1,0.5,1.0,2.0\}$; and for $f^{(\text{BW})}_{n}(t)$ we evaluate $\alpha\in\{2,5,10,15,20\}$. Although Powerformer introduces two more hyperparameters over PatchTST, we observe in Tables S5, S6, S7, and S8 that many masks are not very sensitive to changes in $\alpha$ within a certain range. For this reason, tuning $\alpha$ will likely be very fast as it does not require much fine-tuning after finding this large range.
We note that in addition to the hyperparameters previously discussed, we try two different input sequence lengths (336, 512). There is no standard measurement rate for general time-series datasets, and in general, their dynamics are quite different and happen on different time scales. Thus, the amount of variation within some number of time steps depends on both the dynamics being measured and the measurement rate. The input sequence length must therefore change to encompass the same amount of information (or variation) for datasets with differing measurement rates and temporal variability.
@@
To reproduce our results, we list our hyperparameters used on each dataset for Powerformer (Table S2) and Transformer (Table S3). For some models, we apply weight decay to the encoder only, leaving the forecasting head without regularization. For each dataset and model, we trained with the following 3 random seeds: 2021, 1776, and 1953.
@@
Table S2: We provide Powerformer architecture and training parameters used for each dataset.
@@
Learning Rate $10^{-3}$ $10^{-4}$ $10^{-5}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$
Weight Decay 1.0 1.0 0.05
@@
Table S3: We provide Transformer architecture and training parameters used for each dataset.
@@
Here, we show Powerformer forecasting visualizations and compare various decay time constants against ground truth and PatcTST (Nie et al., 2023). We also highlight plots where both Powerformer and PatchTST struggle. In these cases, Powerformer generally outperforms by better capturing faster-moving features, non-periodic features, and is better able to alter periodic predictions. We believe this is due to Powerformer’s ability to focus on and amplify local high-frequency features while dampening or removing long-time high-frequency components. RBCA achieves this through the power-law masks $f^{(\text{PL})}(t)$ and $f^{(\text{SPL})}(t)$, which amplify local correlations and dampen long-time correlations.
@@
Figure S9: We show forecasting results on the Electricity dataset for Powerformer with $f^{(\text{PL})}(t)$ and PatchTST. The colored lines represent PatchTST and different $f^{(\text{PL})}(t)$ decay constants. For these forecasts, the sequence length is 512 and the prediction length is 336.
@@
Figure S10: We show forecasting results on the ETTm1 dataset for Powerformer with $f^{(\text{PL})}(t)$ and PatchTST. The colored lines represent PatchTST and different $f^{(\text{PL})}(t)$ decay constants. For these forecasts, the sequence length is 512 and the prediction length is 336.
@@
Figure S11: We show forecasting results on the Weather dataset for Powerformer with $f^{(\text{PL})}(t)$ and PatchTST. The colored lines represent PatchTST and different $f^{(\text{PL})}(t)$ decay constants. For these forecasts, the sequence length is 512 and the prediction length is 336.
S5 EXPERIMENTS
S5.1 Flip Flop Experiment Details
Figure S12: The flip-flip OOD evaluation for various recency biases and MHA. The transparent lines are the accuracies, while the solid lines are an exponential moving average.
We generate the flip-flop dataset by producing a string of 512 characters, alternating between instructional characters $c\in\{w,r,i\}$ and numbers $n\int\{0,0\}$ (Press et al., 2022). The character commands dictate the model to do the following: write ($w$) the following number to memory (replacing the previous one), ignore ($i$) the following number and do not commit it to memory, and read ($r$) the number saved in memory. The first character of the string is a $w$, the last character is $r$, and the character following $r$ is space holder ‘-’ that is meant to be replaced by the predicted number. During training, there can be multiple $r$ commands where the most recently written number is predicted. We also generate and test the models on an out-of-distribution (OOD) dataset, where there are many ignore ($i$) commands between the last write and the last read commands. When evaluating the OOD task, we only record results from the last read.
Due to the recurrent neural network (RNNs) architecture, these write-ignore-read capabilities are built in; however, this is not the case for Transformers (Vaswani et al., 2017). To test this, we use a 2-layer Transformer with embedding size 512, max sequence length of 512, absolute positional embeddings, and 4 attention heads. In our experiments, we either varied or removed the causal mask. We trained each model with a learning rate of $10^{-}5$ for 11000 epochs with a batch size of 32 and a weight decay scale of 0.1.
The OOD evaluation as a function of training, Fig. S12, show that the power-law recency biases outperform. In general, the power-law bias $f^{(\text{PL})}(t)$ consistently outperforms regular MHA and other recency biases. Interestingly, the ALiBi bias $f^{(\text{E})}(t)$ does significantly worse than MHA even though it is introducing a recency bias. The similarity power-law $f^{(\text{SPL})}(t)$ performs similarly to MHA, likely because this recency bias is relatively flat.
S5.2 Time-Series Experiment Details
@@
All experiments are evaluated 3 times with different initialization (S3.4). For each dataset, we select a single input sequence based on the best-performing input sequence length across all prediction lengths. We choose a single input sequence length for each dataset because the strength of the correlations between time steps is independent of the prediction length, but the required input sequence length to capture the information necessary to forecast is unique for each dataset. For each prediction length, we treat the attention mask as a hyperparameter and select the best mask and time constant pair. We select attention masks for each prediction length because the forecasting timescale will likely emphasize short or long-term correlations for shorter or longer forecasting windows. We make our selections based on the MSE error. Section S3.4 provides more details on the hyperparameter tuning and the hyperparameters used in this work.
S5.3 Powerformer
@@
We evaluate the weight power-law ($f^{(\text{PL})}(t)$) and similarity power-law ($f^{(\text{SPL})}(t)$) filters on all 7 datasets. For $f^{(\text{PL})}(t)$ we evaluate $\alpha\in\{0.1,0.25,0.5,0.75,1.0\}$ and for $f^{(\text{SPL})}(t)$ we evaluate $\alpha\in\{0.1,0.5,1.0,2.0\}$. Similar to PatchTST (Nie et al., 2023) we evaluate the power-law filters with two input lengths: 336 and 512.
@@
We train a standard encoder-decoder Transformer architecture both with and without RBCA. Generic Transformer models often incorporate causality through masking. To account for this, our Transformer benchmark uses masking for the decoder MHA but not for the encoder MHA nor the decoder cross-attention. We investigate RBCA’s effects by replacing the encoder and decoder self-attention, but not the decoder cross-attention as it is already causal. Section S3.1 provides more architectural details.
@@
We train Transformer models, with and without RBCA, using the weight power-law ($f^{(\text{PL})}(t)$) filter on all 7 datasets. We explore the following decay constants: $\alpha\in[0.1,0.25,0.5,0.75,1.0]$ and provide further experimental detail in sections 4 and S5.2.
S5.5 Baseline Selection
@@
We compare Powerformer against multiple Transformer-based state-of-the-art methods. We populate Table 1 with values taken from the models’ respective papers. All models except TOTEM (Talukder et al., 2024) and iTransformer (Liu et al., 2024) treat the input length as a hyperparameter and report the best results. PatchTST (Nie et al., 2023) provides results for two input sequence lengths (336 and 512). In Table 1 we report the best result between these input sequence lengths for each dataset based on the MSE. We employ the same input sequence length selection for Powerformer.
S5.6 Computing Resources
These experiments were run on NVIDIA A40 and NVIDIA TITAN RTX GPUs. Compute time and resources are highly variable between datasets; the smallest (ETTh1) takes roughly 20 minutes on a single A40, while the largest (Traffic) takes roughly 15 hours on 4 A40s.
S6 TRANSFORMER RESULTS
@@
Here, we present the full results from evaluating Transformer with RBCA. The architecture is described in Appendix S3.1 and the experimental details are provided in Appendices S5.2 and S5.4.
@@
Table S4 provides the aggregate Transformer performance without RBCA and with RBCA for varying decay times ($\alpha$). RBCA decisively outperforms MHA on 4 or the 7 datasets and ties with the number of best results on ETTh2. Looking further into ETTh2 we see that when RBCA outperforms MHA it does so by much larger margins (16%) than when it underperforms (2%). Moreover, we observe the same larger margins of overperformance across all of the datasets. This indicates that when the natural pairwise distribution is found we can see significant improvements in MSE and MAE, but the effects of imposing the wrong pairwise distribution are much less significant.
@@
In Figs. S13-S16 we present the attention score and weight distributions with and without applying the local-causal mask. The distributions can have significant deviations between datasets. In general, the RBCA decoder self-attention shows the smallest deviation from the MHA base case. We note that the MHA base case already has a causal mask. The last RBCA encoder self-attention layer often shows the largest difference between RBCA and MHA distributions. This makes sense since the MHA encoder attention is not causal. We do not alter the decoder cross-attention, but we still observe shifts in the distribution, which differ between datasets. Interestingly, we do not observe the same bimodal distribution in attention weights as we do for Powerformer. This may be due to the Transformer’s encoder-decoder structure, whereas Powerformer only has an encoder.
@@
Table S4: We provide the forecasting MSE and MAE on the test sets for Transformer with MHA (base case) and for Transformer with RBCA and the weight power-law mask $f^{(\text{PL})}(t)$. The top row indicates the model or decay constant $(\alpha)$, the bold numbers are the best (lowest) performance, and the underlined numbers are the second best.
@@
Figure S13: We present the Transformer attention score (inset) and weight distributions on the Weather dataset with a forecasting length of 96. The dotted line represents the reference MHA Transformer results, the dashed-dotted line represents RBCA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,R})}$, and the solid lines represent RBCA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,R})}$.
@@
Figure S14: We present the Transformer attention score (inset) and weight distributions on the Weather dataset with a forecasting length of 720. The dotted line represents the reference MHA Transformer results, the dashed-dotted line represents RBCA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,R})}$, and the solid lines represent RBCA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,R})}$.
@@
Figure S15: We present the Transformer attention score (inset) and weight distributions on the Electricity dataset with a forecasting length of 96. The dotted line represents the reference MHA Transformer results, the dashed-dotted line represents RBCA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,R})}$, and the solid lines represent RBCA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,R})}$.
@@
Figure S16: We present the Transformer attention score (inset) and weight distributions on the Electricity dataset with a forecasting length of 720. The dotted line represents the reference MHA Transformer results, the dashed-dotted line represents RBCA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,R})}$, and the solid lines represent RBCA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,R})}$.
S7 POWERFORMER RESULTS
@@
We present the full Powerformer results on all the evaluated masks: $f^{(\text{PL})}(t)$, $f^{(\text{SPL})}(t)$, $f^{(\text{BW})}_{1}(t)$, and $f^{(\text{BW})}_{2}(t)$. Architecture details can be found in Appendix S3.3, while Appendices S5.2 and S5.3 outline the experimental details.
S7.1 Weight Power-Law Mask
@@
Table S5 presents the aggregate MSE and MAE results for weight power-law mask ($f^{(\text{PL})}(t)$) with varying time decays ($\alpha$), forecasting lengths, and input sequence lengths. The 512 input sequence generally outperforms the shorter 336 input length. As expected, the best performing $\alpha$ varies between datasets, but it also varies as a function of forecast length and input sequence length. Given that some of the best-performing decay lengths are $\alpha=1$ with further experiments one may improve upon these values. However, our $f^{(\text{SPL})}(t)$ mask includes faster-decaying weights, as shown in Fig. 2.
@@
We compare all datasets’ attention score and weight distributions both with MHA and RBCA in Figs S17-S20. We observe similar bimodal distributions for all datasets and note wider $\mathbf{C}^{(\text{C,R})}$ distributions for ETTm1, ETTm2, and Weather.
@@
Table S5: We compare Powerformer’s performance with the weight power law mask $f^{(\text{PL})}(t)$ on standard time-series datasets for varying decay lengths. The best results are bolded and second best are underlined.